A woman is driving 22 miles to work and dropping her daughter at daycare. 1st part of trip is 6 miles less than 2nd part. How long is each part? Create 2 equations in two variables to solve.
Answer by alicealc(293)

1st part = x
2nd part = y
x + y = 22
x = y - 6
substitute the second equation into the first equation:
y - 6 + y = 22
2y - 6 = 22
2y = 22 + 6
2y = 28
y = 28/2 = 14
x = y - 6 = 14 - 6 = 8
so, the 1st part is 8 miles long, and the 2nd part is 14 miles long